Find the vertex for the parabola given below.

f(x) = 3x² + 12x + 16

Answer by Alan3354(69443)

Here's a similar problem:
------------
The vertex of y = -3x²- 6x - 9 lies in which quadrant?
----------
vertex at x = -b/2a = 6/-6 = -1
y(-1) = -3 + 6 - 9 = -6
--> (-1,-6) in Q3